A $5000 investment at an annual simple interest rate of 3.9% earned as much interest after one year as another investment in an account that earned 6.5% annual simple interest. How much was invested at 6.5%?
Answer by jim_thompson5910(35256)

The $5,000 invested at 3.9% simple interest gets you 5000*0.039*1 = 195 dollars in interest alone.

Some other amount, call it x dollars, is invested at 6.5% simple interest for a year. This x dollar investment at 6.5% simple interest will earn 195 dollars in interest alone because the problem states that the two investments earned the same amount of interest.

So

I = Prt

195 = x*0.065*1

195 = 0.065x

0.065x = 195

x = 195/0.065

x = 3000

Which means that 3000 dollars was invested at 6.5% simple interest.